# Diffusion-based Dynamic Contract for Federated AI Agent Construction in Mobile Metaverses

### v1

@@ III-C Energy Consumption of Edge Servers
III-D Utilities of Edge Servers and Cloud Server
@@
where $\alpha$ is a scaling factor that affects the overall magnitude of economic benefits [22], and $E_{C}$ is influenced by several factors [32], including global AI agent integration latency, both CPU and graphics processing unit resources on the cloud, and the average number of instructions in integrating agent modules.
Due to high energy consumption, certain ESs may not always be willing to participate in federated AI agent construction. Thus, ESs have different levels of WTP for federated AI agent construction [16]. The WTP levels of ESs are generally unknown to the cloud server and may change over multiple training rounds [16], which results in dynamic information asymmetry between the cloud server and ESs. Hence, it is necessary to develop a dynamic incentive scheme for ESs to participate in federated AI agent construction sustainably.
@@ IV Dynamic Contract Design
In this section, we propose a two-period dynamic contract model to motivate ESs to continuously participate in federated AI agent construction under dynamic information asymmetry. Note that the proposed dynamic contract model possesses the inherent correlation between two periods within the whole agent construction time and can be easily extended to accommodate multiple periods recursively [16, 22].
@@ IV-A Two-Period Dynamic Contract Formulation
Due to information asymmetry, the cloud server only knows the historical distribution of WTP levels of ESs through previous participation records rather than the precise WTP level of each ES [16, 19]. Thus, the cloud server can use data mining tools to divide ESs into a set $\Theta=\{\theta_{k}|\theta_{k}\in\mathbb{N}_{+},\>k\in\mathcal{K}=\{1,\ldots,K,\>K\leq N\}\}$ of $K$ types based on the historical distribution of WTP levels [16, 19]. Considering the certain correlations in WTP across periods[22, 16], in the first period, the $k$-th type of ESs can be denoted as $\theta_{k}^{1},\>k\in\mathcal{K}$, and in the second period, the $k$-th type of ESs is denoted as $\theta_{j}^{2}(\theta_{k}^{1}),\>j\in\mathcal{K}$, which is an expectation correlated to $\theta_{k}^{1}$. Note that these two types are drawn independently from $\Theta$. Without loss of generality, the types in the first and second periods are sorted in non-decreasing order as $\theta_{1}^{1}\leq\cdots\leq\theta_{k}^{1}\leq\cdots\leq\theta_{K}^{1}$ and $\theta_{1}^{2}(\theta_{k}^{1})\leq\cdots\leq\theta_{j}^{2}(\theta_{k}^{1})\leq\cdots\leq\theta_{K}^{2}(\theta_{k}^{1})$, respectively. We define $p_{k}^{t}$ as the probability that ESs belong to type $\theta_{k}^{t}$ during period $t$, where $\sum_{k=1}^{K}p_{k}^{t}=1,\>t\in\{1,2\}$.
To incentivize ESs to participate in federated AI agent construction, the cloud server issues a two-period dynamic contract $\Omega(\theta_{k}^{1},\theta_{j}^{2}(\theta_{k}^{1}))$ at the beginning of the first period before the agent module creation begins. The contract item consists of WTP along with the corresponding reward. In the first period, the contract item designed for type-$k$ ESs is represented as $\{T_{k}^{1}(\theta_{k}^{1}),R_{k}^{1}(\theta_{k}^{1})\}$, and in the second period, the contract item designed for the same type-$k$ ESs is represented as $\{T_{j}^{2}(\theta_{j}^{2}(\theta_{k}^{1})),R_{j}^{2}(\theta_{j}^{2}(\theta_{k}^{1}))\},\>k,j\in\mathcal{K}$. Thus, the formulation of the two-period dynamic contract $\Omega$ is expressed as
@@
where $\beta\in(0,1]$ represents a discount factor for the utility in period 2. The second term in (10) represents the expected discounted utility across all $K$ types that users are characterized with certain probabilities in period 2, and a higher $\beta$ indicates a greater valuation of the future utility relative to the current utility [22, 16]. Since WTP levels across two periods are positively correlated [16], we have $p_{j}^{2}(T_{k}^{1})\neq p_{j}^{2}(T^{1}_{k^{\prime}})$ and $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))\neq\sum_{j=1}^{K}p_{j}^{2}(T^{1}_{k^{\prime}})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}})),\>k\neq k^{\prime}$, with proofs presented in Lemma 7.
@@
Figure 2: The workflow of the proposed two-period dynamic contract for agent module construction. To minimize disruptions to the construction progress, the proposed contract remains valid across both periods without requiring redefinition and retransmission to the ESs in period 2.
@@
  • Stage III. Contract realization of period 1: At the end of period 1, the cloud server integrates the created agent modules. If the integration is successful, the cloud server will offer the rewards $R_{k}^{1}(\theta_{k}^{1})$ specified in the contract to ESs. Otherwise, the ESs will not obtain the rewards.
@@ IV-B Second-Period Contract Design
IR constraints in (13) motivate ESs to participate in federated AI agent construction by guaranteeing that their utilities are non-negative if they select suitable contract items. IC constraints in (14) encourage ESs to reveal their actual type, avoiding the risk that ESs misrepresent their actual types to cheat higher rewards. In the following, we derive the necessary conditions of the second-period contract.
@@
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1}).$ (15)
We transform the above inequality into
$\theta_{i}^{2}[R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1})]\geq c[T_{i}^{2}(\theta_{k}^{1})-T_{j}^{2}(\theta_{k}^{1})].$ (16)
Since $T_{i}^{2}(\theta_{k}^{1})\geq T_{j}^{2}(\theta_{k}^{1})$, we have $\theta_{i}^{2}[R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1})]\geq 0$, then $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$. Similarly, given that $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$, we have $T_{i}^{2}(\theta_{k}^{1})\geq T_{j}^{2}(\theta_{k}^{1})$. Thus, the proof of Lemma 1 is completed. ∎
@@
Based on the IC constraints in (14), we have
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1}),$ (17)
$\theta_{j}^{2}R_{j}^{2}(\theta_{k}^{1})-cT_{j}^{2}(\theta_{k}^{1})\geq\theta_{j}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1}).$ (18)
By combining (17) and (18), we have
$(\theta_{i}^{2}-\theta_{j}^{2})(R_{i}^{2}(\theta_{k}^{1})-R_{j}^{2}(\theta_{k}^{1}))\geq 0.$ (19)
Since $\theta_{i}^{2}(\theta_{k}^{1})\geq\theta_{j}^{2}(\theta_{k}^{1})$, we have $R_{i}^{2}(\theta_{k}^{1})\geq R_{j}^{2}(\theta_{k}^{1})$. Thus, the proof of Lemma 2 is completed. ∎
Lemma 1 ensures that ESs can receive higher rewards with more training rounds for improving the model quality of agent modules. Lemma 2 indicates that ESs with higher WTP can obtain higher rewards. Based on Lemmas 1 and 2, we can obtain the monotonicity property of the second-period contract as follows:
@@
Since it is exceedingly challenging to make the contract solution under $K$ number of IR constraints and $K(K-1)$ number of IC constraints, we reduce the numbers of IR and IC constraints to derive a tractable set of constraints to solve the second-period contract.
@@
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})&-E\geq\theta_{i}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\\ &\geq\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\geq 0.\end{split}$ (21)
From (21), we can observe that if the type-$1$ user with the lowest WTP has a non-negative utility, it implies that the other IR constraints automatically hold. Additionally, for $\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E\geq 0$, the cloud server would reduce the reward $R_{1}^{2}(\theta_{k}^{1})$ as much as possible until $\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E=0$ [19], thus maximizing its utility. Thus, the proof of Lemma 4 is completed. ∎
@@
$\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1}).$ (24)
$\theta_{i-1}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i-1}^{2}R_{i-2}^{2}(\theta_{k}^{1})-cT_{i-2}^{2}(\theta_{k}^{1}),$ (25)
where (25) can be transformed into
$\theta_{i-1}^{2}(R_{i-1}^{2}(\theta_{k}^{1})-R_{i-2}^{2}(\theta_{k}^{1}))\geq c(T_{i-1}^{2}(\theta_{k}^{1})-T_{i-2}^{2}(\theta_{k}^{1})).$ (26)
Based on the ascending order of types, we can further obtain
$\theta_{i}^{2}(R_{i-1}^{2}(\theta_{k}^{1})-R_{i-2}^{2}(\theta_{k}^{1}))\geq c(T_{i-1}^{2}(\theta_{k}^{1})-T_{i-2}^{2}(\theta_{k}^{1})),$ (27)
which can be transformed into
$\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i}^{2}R_{i-2}^{2}(\theta_{k}^{1})-cT_{i-2}^{2}(\theta_{k}^{1}).$ (28)
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq&\theta_{i}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\\ &\cdots\geq\theta_{i}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1}).\end{split}$ (29)
$\begin{split}\theta_{i}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1})\geq&\theta_{i}^{2}R_{i+1}^{2}(\theta_{k}^{1})-cT_{i+1}^{2}(\theta_{k}^{1})\geq\\ &\cdots\geq\theta_{i}^{2}R_{K}^{2}(\theta_{k}^{1})-cT_{K}^{2}(\theta_{k}^{1}).\end{split}$ (30)
Thus, the proof of Lemma 5 is completed. ∎
@@
(Reduce IC Constraints in Period 2): The $K(K-1)$ number of IC constraints in period 2 can be reduced to $(K-1)$ number of IC constraints, i.e.,
@@
$\theta_{i}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1}))\geq\theta_{i-1}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1})),$ (32)
which (32) can be further transformed into
$c(T_{i}^{2}(\theta_{k}^{1})-T_{i-1}^{2}(\theta_{k}^{1}))\geq\theta_{i-1}^{2}(R_{i}^{2}(\theta_{k}^{1})-R_{i-1}^{2}(\theta_{k}^{1})),$ (33)
and we have
$\theta_{i-1}^{2}R_{i-1}^{2}(\theta_{k}^{1})-cT_{i-1}^{2}(\theta_{k}^{1})\geq\theta_{i-1}^{2}R_{i}^{2}(\theta_{k}^{1})-cT_{i}^{2}(\theta_{k}^{1}).$ (34)
Therefore, the LUIC constraints remain valid. The proof of Lemma 6 is completed. ∎
Based on Lemmas 4 and 6, the $K$ number of IR constraints (13) and $K(K-1)$ number of IC constraints (14) can be reduced to one IR constraint (20) and $(K-1)$ number of IC constraints (31). We derive the sufficient and necessary conditions of the feasible contract in period 2 as follows:
@@
$3)\quad\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E=0$
$4)\quad\theta_{i}^{2}R_{i}^{2}-cT_{i}^{2}-E=\theta_{i}^{2}R_{i-1}^{2}-cT_{i-1}^{2}-E,\>\forall i\geq 2$
@@
Please refer to Appendix A. ∎
@@ IV-C First-Period Contract Design
For the feasible contract design in period 1, the IR constraints are similar to that in Definition 1, which is given by
@@
From the IIC constraints in (37), we know that ESs will obtain higher utilities by truthfully declaring their WTP levels. When the first-period and second-period types are mutually independent, we have $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))=\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))$, where $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))$ represents the discounted expected utility of the type-$k$ ESs that declare their types are $\theta^{1}_{k^{\prime}}$ in period 1, and the IIC constraints can be simplified as the IC constraints of the standard static contract. In the following, we derive the necessary conditions of the feasible contract in period 1.
@@
Please refer to Appendix B. ∎
@@
Please refer to Appendix C. ∎
Lemma 7 indicates that the ESs with higher WTP levels will perform more training rounds for agent module creation in both period 1 and period 2, even if they have the same WTP level in period 2. Lemma 8 represents the monotonicity of first-period rewards and training rounds, indicating that the ESs with higher WTP levels will receive higher rewards from the cloud server in period 1.
In the following, we reduce the IR and IIC constraints in period 1, which is similar to period 2.
@@
(Reduce IIC Constraints in Period 1): Similar to (31), the $K(K-1)$ number of IIC constraints can be reduced into the $(K-1)$ number of local downward IIC constraints, which is given by

### v2

@@ III-C Energy Consumption of Edge Servers
III-D Utilities of Edge Servers and the Cloud Server
@@
where $\alpha$ represents a scaling factor that affects the overall magnitude of economic benefits [7], and $E_{C}$ is influenced by several factors [35], including the latency of global AI agent integration, both CPU and Graphics Processing Unit (GPU) resources on the cloud, and the average number of instructions in integrating agent modules.
Due to high workload and energy consumption, ESs may not always be willing to participate in federated AI agent construction, indicating that ESs have different levels of WTP for federated AI agent construction [14]. The WTP levels of ESs are generally unknown to the cloud server and may change over multiple training rounds [14], which results in dynamic information asymmetry between the cloud server and ESs. Hence, it is necessary to develop a dynamic incentive scheme for ESs to participate in federated AI agent construction.
@@ IV Dynamic Contract Design
In this section, we propose a two-period dynamic contract model to motivate ESs to continuously participate in federated AI agent construction under dynamic information asymmetry. Note that the proposed dynamic contract model possesses the inherent correlation between two periods within the whole agent construction time and can be feasibly extended to accommodate multiple periods recursively [14, 7]. In practice, the two-period dynamic contract design is efficient to balance incentive effectiveness and computing complexity [14].
@@ IV-A Two-Period Dynamic Contract Formulation
Due to information asymmetry, the cloud server only knows the historical distribution of WTP levels of ESs through previous participation records rather than the precise WTP level of each ES [14, 31]. Thus, the cloud server can use data mining tools to divide ESs into a set $\Theta=\{\theta_{k}|\theta_{k}\in\mathbb{N}_{+},\>k\in\mathcal{K}=\{1,\ldots,K,\>K\leq N\}\}$ of $K$ types based on the historical distribution of WTP levels [14, 31]. Considering the certain correlations in WTP across periods [7, 14], in the first period, the $k$-th type of ESs can be denoted as $\theta_{k}^{1},\>k\in\mathcal{K}$, and in the second period, the $k$-th type of ESs is denoted as $\theta_{j}^{2}(\theta_{k}^{1}),\>j\in\mathcal{K}$, which is an expectation correlated to $\theta_{k}^{1}$. Note that these two types are drawn independently from $\Theta$. Without loss of generality, the types in the first and second periods are sorted in non-decreasing order as $\theta_{1}^{1}\leq\cdots\leq\theta_{k}^{1}\leq\cdots\leq\theta_{K}^{1}$ and $\theta_{1}^{2}(\theta_{k}^{1})\leq\cdots\leq\theta_{j}^{2}(\theta_{k}^{1})\leq\cdots\leq\theta_{K}^{2}(\theta_{k}^{1})$, respectively [14, 7]. We define $p_{k}^{t}$ as the probability that ESs belong to type $\theta_{k}^{t}$ during period $t$, where $\sum_{k=1}^{K}p_{k}^{t}=1,\>t\in\{1,2\}$.
To incentivize ESs to continuously participate in federated AI agent construction, the cloud server issues a two-period dynamic contract $\Omega(\theta_{k}^{1},\theta_{j}^{2}(\theta_{k}^{1}))$ at the beginning of the first period before the agent module creation starts. The contract item consists of WTP along with the corresponding reward. In the first period, the contract item designed for type-$k$ ESs is represented as $\{T_{k}^{1}(\theta_{k}^{1}),R_{k}^{1}(\theta_{k}^{1})\}$, and in the second period, the contract item designed for the same type-$k$ ESs is represented as $\{T_{j}^{2}(\theta_{j}^{2}(\theta_{k}^{1})),R_{j}^{2}(\theta_{j}^{2}(\theta_{k}^{1}))\},\>k,j\in\mathcal{K}$. Thus, the formulation of the two-period dynamic contract $\Omega$ is expressed as
@@
where $\beta\in(0,1]$ and $p_{j}^{2}(T_{k}^{1})$ represent a discount factor for the utility and the probability that type-$k$ ESs belong to type $\theta_{j}^{2}$ during period 2, respectively. The second term in (10) represents the expected discounted utility across all $K$ types that ESs are characterized with certain probabilities in period 2, which is designed to capture the long-term benefit of ESs participating in the federated AI agent construction. A larger $\beta$ indicates a greater valuation of the future utility relative to the current utility [7, 14]. Since WTP levels across two periods are positively correlated [14], we have $p_{j}^{2}(T_{k}^{1})\neq p_{j}^{2}(T^{1}_{k^{\prime}})$ and $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))\neq\sum_{j=1}^{K}p_{j}^{2}(T^{1}_{k^{\prime}})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}})),\>k\neq k^{\prime}$, with the proof presented in Lemma 7 in this paper.
@@
Figure 2: The workflow of the proposed two-period dynamic contract for agent module construction. To minimize disruptions to the construction progress, the proposed dynamic contract remains valid across both periods without requiring redefinition and retransmission to the ESs in period 2.
@@
  • Stage III. Contract realization of period 1: At the end of period 1, the cloud server integrates the created agent modules. If the integration is successful, the cloud server will offer the rewards $R_{k}^{1}(\theta_{k}^{1})$ specified in the contract to ESs. Otherwise, the ESs that fail to fulfill the contractual requirements will not obtain the corresponding rewards.
@@ IV-B Second-Period Contract Design
IR constraints in (13) motivate ESs to participate in federated AI agent construction by guaranteeing that their utilities are non-negative if they select suitable contract items. IC constraints in (14) encourage ESs to reveal their actual type, avoiding the risk that ESs misrepresent their actual types to cheat for higher rewards. In the following, we derive the necessary conditions of the second-period contract.
@@
Please refer to Appendix A. ∎
@@
Please refer to Appendix B. ∎
Lemma 1 ensures that ESs can receive higher rewards with more training rounds for improving the model quality of agent modules. Lemma 2 indicates that ESs with a higher WTP level can obtain higher rewards. According to Lemmas 1 and 2, we can prove the monotonicity property of the second-period contract as follows:
@@
Since it is exceedingly challenging to make the dynamic contract solution under $K$ number of IR constraints and K(K1)K(K-1) number of IC constraints, we reduce the number of IR and IC constraints to derive a tractable set of constraints to solve the second-period contract.
@@
Please refer to Appendix C. ∎
@@
Please refer to Appendix D. ∎
@@
(Reduce IC Constraints in Period 2): The K(K1)K(K-1) number of IC constraints in period 2 can be reduced to (K1)(K-1) number of IC constraints, i.e.,
@@
Please refer to Appendix E. ∎
Based on Lemmas 4 and 6, the $K$ number of IR constraints (13) and K(K1)K(K-1) number of IC constraints (14) can be reduced to one IR constraint (15) and (K1)(K-1) number of IC constraints (18). We derive the sufficient and necessary conditions of the feasible contract in period 2 as follows:
@@
$3)\quad\theta_{1}^{2}R_{1}^{2}(\theta_{k}^{1})-cT_{1}^{2}(\theta_{k}^{1})-E=0.$
$4)\quad\theta_{i}^{2}R_{i}^{2}-cT_{i}^{2}-E=\theta_{i}^{2}R_{i-1}^{2}-cT_{i-1}^{2}-E,\>\forall i\geq 2.$
@@
Please refer to Appendix F. ∎
@@ IV-C First-Period Contract Design
For the feasible contract design in period 1, the IR constraints are similar to those in Definition 1, as given by
@@
From the IIC constraints in (21), we know that ESs will obtain higher utilities by truthfully declaring their WTP levels. When the first-period and second-period types are mutually independent, we have $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))=\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))$, where $\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))$ represents the discounted expected utility of the type-$k$ ESs that declare their types are $\theta^{1}_{k^{\prime}}$ in period 1, and the IIC constraints can be simplified as the IC constraints of the standard static contract. In the following, we derive the necessary conditions for the feasible contract in period 1.
@@
Please refer to Appendix G. ∎
@@
Please refer to Appendix H. ∎
Lemma 7 indicates that the ESs with higher WTP levels will perform more training rounds for agent module creation in both period 1 and period 2, even if they have the same WTP level in period 2. Lemma 8 represents the monotonicity of first-period rewards and training rounds, indicating that the ESs with higher WTP levels will receive higher rewards from the cloud server in period 1. Next, we reduce the IR and IIC constraints in period 1, which is similar to period 2.
@@
(Reduce IIC Constraints in Period 1): Similar to (18), the K(K1)K(K-1) number of IIC constraints can be reduced into the (K1)(K-1) number of local downward IIC constraints, which is given by